OpenStudy (anonymous):

What is the oblique asymptote of the function f(x) = the quantity of x squared plus x minus 2, all over x plus 1? @FriedRice

OpenStudy (dan815):

x^2+x-2 --------- x+1

OpenStudy (dan815):

= x(x+1) - 2 ------ ------ x+1 x+1 =x - 2/(x+1)

OpenStudy (anonymous):

What is the oblique asymptote of the function

OpenStudy (dan815):

what happens to function for x=-1 for that 2nd term

OpenStudy (anonymous):

oh I get it now thanks